# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v6

@@ 5.4 Other Use Cases
To demonstrate that SMD is not limited to the three use cases we primarily focus on, we discuss two other use cases.
@@
Online DRAM Error Profiling. SMD enables implementing a maintenance mechanism for online profiling of DRAM errors that may occur when operating with increased refresh period and reduced DRAM timing parameters [34, 17, 106, 107, 108, 1, 109, 110, 86, 38, 15, 111]. Such a maintenance mechanism can be used to explore how DRAM cell characteristics vary throughout the DRAM chip. To profile a DRAM region, the maintenance mechanism can use SMD to lock this region and prevent the MC from accessing it while the region is being profiled. Because the profiling operation can cause bit flips, the maintenance mechanism should temporarily buffer the original data that is stored in the profiled region. For example, for profiling errors that occur at reduced tRCD, the maintenance mechanism should first safely copy the data stored in a row to another storage space (e.g., an unused row or an SRAM buffer), before attempting to activate the row with reduced tRCD. The profiler mechanism can internally share the profiling results with other maintenance mechanisms (e.g., for skipping refreshes to rows with high retention time) and mechanisms for improving DRAM access latency and energy efficiency.
@@
Processing in/near Memory. In the presence of an in-DRAM processing engine, SMD can help resolve access conflicts between the in-DRAM processing engine and the MC. To do so, SMD can treat the in-DRAM processing engine as a maintenance mechanism. The in-DRAM processing engine can use SMD to lock a DRAM region that it will operate on. Because SMD does not allow the MC to activate a row in a locked region, only the in-DRAM processing engine will have access to the locked region until it completes the processing and releases the region.
@@ 6 Experimental Methodology
We extend Ramulator [82, 81] to implement and evaluate the three SMD maintenance mechanisms (SMD-FR, SMD-DRP, and SMD-MS) that we describe in section 5. We use DRAMPower [112, 113] to evaluate DRAM energy consumption. We use Ramulator in CPU-trace driven mode executing traces of representative sections of our workloads collected with a custom Pintool [114]. We warm-up the caches by fast-forwarding 100 million (M) instructions. We simulate each representative trace for 500M instructions (for multi-core simulations, until each core executes at least 500M instructions).
We use the system configuration provided in Table 1 in our evaluations. Although our evaluation is based on DDR4 DRAM, the modifications required to enable SMD can be adopted in other DRAM standards, as we explain in section 7.4.
@@
| Memory Controller | 64-entry read/write request queue, FR-FCFS-Cap [115] |
@@
DDR4 Baseline. The baseline system 1) uses per-rank refresh111111The DDR4 standard does not support per-bank refresh. and 2) does not perform ECC-scrubbing.
@@
Comparison Points. We compare SMD to memory controller/DRAM co-design techniques. First, DARP [16] intelligently schedules per-bank refresh commands to idle banks while the memory controller is in write mode to reduce delays imposed by refresh operations on memory demand requests (e.g., load instructions executed by the processor). Second, DSARP [16] implements DARP and modifies DRAM chips and the interface to perform a refresh operation in one subarray and a memory accesses in another subarray and thereby hide the latency of refresh operations and reduce delays they impose on memory requests. We modify DARP and DSARP such that they consider victim row refresh operations issued by Deterministic RowHammer Protection (section 5.2) and attempt to reduce delays these operations impose on memory requests.
@@
Configuring Maintenance Mechanisms. SMD-FR refreshes a DRAM row every . Based on [17], for SMD-VR, we conservatively assume 0.1% of rows in each bank need to be refreshed every while the rest retain their data correctly for and more. SMD-VR uses a -bit Bloom Filter with hash functions. SMD-PRP and SMD-PRP+ refresh the victims of an activated row with a high probability, i.e., . SMD-DRP refreshes the victims before the aggressor is activated (512) times during a tREFW (). SMD-MS operates with an aggressive 5-minute scrubbing period.
@@
Evaluated System Configurations. SMD-FR-1LR implements a single lock region in a bank, i.e., it cannot concurrently perform a refresh operation and a memory access. DSARP/SMD-FR can concurrently perform a refresh in a subarray/lock region and a memory access in another. MC-Combined-DARP, MC-Combined-DSARP, and SMD-Combined implement the memory-controller-(MC) and SMD-based versions of Fixed Rate Refresh (section 5.1), Deterministic RowHammer Protection (section 5.2), and Memory Scrubbing (section 5.3). MC-based implementations of SMD maintenance mechanisms are functionally equivalent to SMD-based ones (e.g., SMD-DRP and MC-DRP are configured to provide the same RowHammer prevention guarantees).
7 Evaluation
We evaluate the performance and energy efficiency of SMD-based maintenance mechanisms.
@@ 7.1 Single-core Performance
Fig. 9 shows the speedup of the 22 highest MPKI single-core workloads for DARP, DSARP, SMD, and a hypothetical No-Refresh configuration that does not perform any maintenance operations, over the Baseline system.
We make three key observations. First, SMD-Combined provides a comparable average speedup of 5.9% to No-Refresh’s 5.0% as SMD-Combined eliminates a large fraction of the maintenance overhead by allowing the memory controller to access lock regions that are not under maintenance. Second, SMD-Combined outperforms MC-Combined-DSARP and MC-Combined-DARP by 0.6% and 3.9%, respectively. While MC-Combined-DSARP can concurrently access main memory and perform a maintenance operation, it needs to issue a DRAM command (e.g., a per bank refresh command for periodic refresh) for each maintenance operation and incur delays for other DRAM commands that the memory controller issues to serve main memory requests. In contrast, SMD-Combined autonomously performs each maintenance operation inside the DRAM chip and does not incur delays for DRAM commands (except those incurred by ACT_NACKs). MC-Combined-DARP cannot concurrently perform a maintenance operation and access main memory, thereby performs worse than both SMD-Combined and MC-Combined-DSARP. We make similar observations on SMD-FR’s, DSARP’s, and DARP’s performance. These observations are in line with prior work [16]. Third, intelligent refresh scheduling (DARP) provides 1.1% average speedup across all workloads by reducing the impact of refresh operations on the latency of memory requests. SMD-FR-1LR incurs 2.9% slowdown over DARP. Although SMD-FR-1LR locks a bank only when there is no active row in a bank, it does not have a global view over memory requests in memory controller read/write queues, unlike DARP, and cannot opportunistically perform refresh operations while the memory controller is in write mode.
@@ 7.2 Multi-core Performance
Fig. 10 shows weighted speedup (normalized to the weighted speedup of the DDR4 baseline) for 60 four-core workloads (20 per memory intensity level). The white circles inside each box (error lines) represent the average (minimum and maximum) weighted speedup across the 20 workloads in the corresponding group.
We make three major observations from Fig. 10. First, SMD-Combined provides 8.6% and 4.1% speedups on average across 4c-high workloads over MC-Combined-DARP and MC-Combined-DSARP, respectively. We attribute these speedups to i) SMD’s maintenance-access parallelization and ii) SMD’s ability to perform maintenance operations autonomously inside the DRAM chip without the memory controller having to issue DRAM commands. Second, SMD-FR-1LR incurs an average 4.5% slowdown on average across 4c-high workloads compared to DARP. While DARP can reduce the overheads of periodic refresh by intelligently scheduling refresh operations, SMD-FR-1LR cannot. We attribute SMD-FR-1LR’s overheads to the relatively high rate of ACT_NACK commands (not shown in the figure) it issues: SMD-FR-1LR issues an ACT_NACK command for every 11.7 activate commands, on average across all 4c-high workloads. Third, we observe that i) the speedup and slowdown trends for the tested mechanisms are similar for the average 4c-medium workload, and ii) the tested mechanisms provide little performance improvement for the average 4c-low workload with very low memory intensity.
@@
We conclude that SMD i) provides substantial system performance benefits by concurrently performing a maintenance operation and a memory access in different lock regions in a bank and ii) does not significantly hurt system performance when SMD performs maintenance at bank granularity (i.e., implements one lock region per DRAM bank).
7.3 SMD’s Energy Consumption

### v7

@@ 5.4 Other Use Cases
To demonstrate that SMD can enable many other use cases, we describe three other use cases.
@@
Online DRAM Error Profiling. SMD enables implementing a maintenance mechanism for online profiling of DRAM errors that may occur when operating with increased refresh period and reduced DRAM timing parameters [37, 20, 187, 188, 189, 1, 190, 191, 161, 41, 17, 156, 6, 4, 192, 193, 115]. Such a maintenance mechanism can be used to exploit the variation in DRAM cell characteristics within a DRAM chip. To profile a DRAM region, the maintenance mechanism can use SMD to lock the region and prevent the MC from accessing it while the region is being profiled. Because the profiling operation can cause bit flips, the maintenance mechanism should temporarily buffer the original data that is stored in the profiled region [187, 41]. For example, for profiling errors that occur at reduced tRCD, the maintenance mechanism should first safely copy the data stored in a row to another storage space (e.g., an unused row or an SRAM buffer), before attempting to activate the row with reduced tRCD. The profiling mechanism can internally share the profiling results with other maintenance mechanisms (e.g., maintenance mechanisms 1) for skipping refreshes to rows with high retention times [41, 17, 20, 37, 161] and 2) changing the aggressiveness of RowHammer protection mechanisms based on per-row vulnerability [115]) and mechanisms for improving DRAM access latency and energy efficiency [188, 187, 1, 194].
@@
Processing in/near Memory. In the presence of an in-DRAM processing engine (as in e.g., [195, 196, 197, 198, 145, 146, 147]) , SMD can help resolve access conflicts between the in-DRAM processing engine and the MC. To do so, SMD can treat the in-DRAM processing engine as a maintenance mechanism. The in-DRAM processing engine can use SMD to lock a DRAM region that it will operate on. Because SMD does not allow the MC to activate a row in a locked region, only the in-DRAM processing engine will have access to the locked region until it completes the processing and releases the region.
Power Management. Modern DRAM chips implement self-refresh and power-down modes to save energy [119, 141]. The whole DRAM chip is inaccessible in power-down mode, such that the DRAM chip does not execute DRAM commands (e.g., ACT, PRE, RD, WR), but internally refreshes DRAM cells (in self-refresh mode) to maintain data integrity. Although a DRAM chip consumes significantly lower energy in self-refresh and power-down modes, a memory request accessing a small piece of data (e.g., 64B of a cache block out of 16GB of a DRAM module) causes the DRAM chip to completely get out of self-refresh or power-down modes and consume significantly larger energy. To make matters worse, such mode switching introduces additional delays for memory requests [141, 119, 199]. To overcome these challenges, SMD can be leveraged to partially power down DRAM chips by powering down rarely accessed memory regions, such that DRAM energy consumption can be reduced for a large portion of the DRAM chip while memory requests targeting frequently accessed memory regions do not experience additional delays. Because SMD does not allow the MC to activate a row in a locked (e.g., powered down) region, only activate commands targeting powered-up regions would proceed without getting ACT_NACK’d. SMD would reject activate commands targeting powered-down memory regions until those regions are powered on.
5.5 Other Maintenance Operation Improvements
SMD can seamlessly enable optimizations for the three maintenance operations we extensively describe (SMD-FR section 5.1, SMD-DRP section 5.2, and SMD-MS 5.3). At a high-level, any optimization idea or technique that is independent from the organization of multiple DRAM chips into a module or a rank (i.e., any idea that can be leveraged by some technique in a single DRAM chip) can be implemented in an SMD chip. Implementing such an idea or technique in an SMD chip does not require further changes to the DRAM interface.
Refresh Overhead Mitigation. A large set of prior work proposes various ideas and techniques that mitigate DRAM refresh overhead (e.g., [19, 20, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]). Many of these ideas and techniques could improve in-DRAM autonomous maintenance operations. One prior technique mitigates DRAM refresh overhead [45] by cleverly leveraging DRAM module organization and rank-level ECC to alleviate the DRAM bandwidth overhead of refresh operations. While SMD does not preclude this technique from being implemented, implementing it as mainly described in [45] would require modifications to the DRAM interface. However, the key idea of [45] could still be leveraged inside an SMD chip by combining 1) on-die ECC [176, 177, 178, 10, 179, 180, 181, 182, 183], which already exists (e.g., in DDR5 [119]) and 2) finer-granularity (DRAM-mat-level) refresh operations (as described in [45]) in a future DRAM design.
Other RowHammer Mitigations. Many prior works provide RowHammer mitigation techniques (e.g., [114, 115, 116, 200, 56, 63, 61, 59, 57, 54, 201, 12, 64, 202, 71, 70, 67, 99, 101, 102, 107]). The key ideas of these works can be implemented using SMD to enable more performance-, area-, and energy-efficient and robust in-DRAM RowHammer protection.
6 Hardware Implementation and Overhead
6.1 DRAM Interface Modifications
An SMD chip needs to send ACT_NACK signals to the memory controller (MC). SMD can 1) repurpose the existing alert_n pin in DDR4 chips [141, 203] or 2) introduce an extra physical pin to transmit ACT_NACK signals.
1) alert_n. alert_n is currently used to inform the MC that the DRAM chip detected a CRC or parity check failure on the issued command, and thus the MC must issue the command again. The alert_n signal can simply be asserted not only on CRC or parity check failure, but also when the MC attempts accessing a row in a locked region. This approach does not require an additional physical pin. However, 1) assigning multiple meanings to an already ambiguously-defined alert_n signal [141] could complicate MC design and 2) alert_n, as currently defined in the standard [141], is an open drain signal that is set and reset slowly.
2) Introducing a new pin. To potentially simplify MC design, SMD can introduce a new pin to transmit ACT_NACK signals. Only one new pin between the MC and all memory devices (e.g., all DRAM chips in a memory channel or a rank) is enough for systems that use rank-based DRAM chip organizations (e.g., DDR5 DIMMs). For example, 96 new pins are needed for a high-end system with 12 memory module channels equipped with 4-rank DDR5 modules in each channel (each DDR5 rank has two memory channels). However, a system of this scale already has >6K total processor pins (e.g., 6,096 in AMD SP5 Socket [204]), and 96 new pins would amount to a relatively small 1.6% increase in pins. Similar to how the alert_n signal works, the module can send the MC a single ACT_NACK when any of the per-chip ACT_NACK signals is asserted.
6.2 DRAM Chip Modifications
DRAM Chip Modifications for SMD. We use CACTI 6.0 [205] to evaluate the hardware overhead of the changes that SMD introduces in a conventional DRAM bank (highlighted in Fig. 2) assuming technology.161616Capitalizing on the latest DRAM technology libraries to implement the changes that SMD introduces would likely provide more accurate area overhead estimations. However, such libraries and tools are proprietary. Lock Region Bitvector (LRB) is a small bitvector that stores a single bit for each lock region in a DRAM bank to indicate whether or not the lock region is under maintenance. In our evaluation, we assume that a bank is divided into lock regions. Therefore, LRB consists of only bits, which are indexed using a 4-bit lock region address. According to our evaluation, an LRB incurs only area overhead per bank. The area overhead of all LRBs in a DRAM chip is only 0.001% of a DRAM chip. The access time of an LRB is , which is only 0.4% of typical row activation latency (tRCD) of .
SMD adds a per-region RA-latch to enable accessing one lock region while another is under maintenance. An RA-latch stores a pre-decoded row address, which is provided by the global row address decoder, and drives the local row decoders where the row address is fully decoded. This design builds on the basic design proposed in SALP [148] and refresh-access parallelization introduced in [19, 149]. According to our evaluation, all RA-latches incur a total area overhead of 1.1% in a typical DRAM chip. An RA-latch has only latency, which is negligible compared to tRCD.
DRAM Chip Modifications for Maintenance Mechanisms. Besides these changes that are the core of the SMD substrate, a particular maintenance mechanism may incur additional area overhead. We evaluate the DRAM chip area overhead of the three maintenance mechanisms presented in section 5.1, section 5.2, and section 5.3. The simple fixed-rate refresh mechanism, SMD-FR, requires only additional area in a DRAM chip. The DRAM chip area overhead of the refresh mechanism is less than 0.1% of a typical DRAM chip. SMD-DRP requires a large Counter Table with 1224 counters per bank for the RowHammer threshold value that we use in our performance evaluation. SMD-DRP requires area, which is 7.0% of a typical DRAM chip.171717Note that the area overhead comes especially from Graphene’s expensive CAM structures [61]. Cheaper RowHammer mitigation mechanisms can be implemented with SMD, as we describe in section 5.5. The control logic of SMD-MS is similar to the control logic of SMD-FR and it requires only additional area excluding the area of the ECC engine, which is already implemented by DRAM chips that support in-DRAM ECC.
6.3 Memory Controller Modifications
We slightly modify the MC’s scheduling mechanism to retry a rejected ACT command as we explain in section 4.3. Upon receiving an ACT_NACK, the MC marks the bank as precharged. An existing MC already implements control circuitry to pick an appropriate request from the request queue and issue the necessary DRAM command based on the DRAM bank state (e.g., ACT to a precharged bank or RD if the corresponding row is already open) by respecting the DRAM timing parameters. The ACT Retry Interval (ARI) is simply a new timing parameter that specifies the minimum time interval for issuing an ACT to a lock region after receiving an ACT_NACK from the same region. Therefore, SMD can be implemented in existing MCs with only slight modifications by leveraging the existing request scheduler (e.g., FR-FCFS [160]).
Tracking Locked Regions. To apply the ARI timing parameter (e.g., to an ACT command targeting a locked region), the memory controller tracks which regions are locked. The memory controller could store the address of the locked region in every DRAM bank (since only one region can be under maintenance at a time in a bank). 1 rank address bit, 4 bank address bits, and 4 lock region address bits (9 bits) are sufficient to track every under-maintenance lock region in 2 ranks and 16 banks (32 unique banks) across all chips on a DRAM module. Therefore, the storage cost for tracking locked regions is only 288 bytes for the evaluated memory channel with a dual-rank x8 memory module.
Address Mapping Schemes. SMD does not require modifications to how a physical address is mapped to a memory-controller-visible DRAM address. In section 8, we evaluate SMD using a mapping scheme that interleaves consecutive cache blocks across channels, columns, ranks, banks, and rows, in that order. We also evaluate SMD using the mapping scheme described in [206] that aims to exploit bank-level parallelism. For this mapping scheme, SMD-FR provides 8.7% speedup over the baseline system, on average across 4c-high workloads. We leave detailed evaluation of SMD with different address mapping schemes for future work.
@@ 7 Experimental Methodology
We extend Ramulator [151, 150] to implement and evaluate the three SMD maintenance mechanisms (SMD-FR, SMD-DRP, and SMD-MS) that we describe in section 5.1, section 5.2, and section 5.3. We use DRAMPower [207, 208] to evaluate DRAM energy consumption. We use Ramulator in CPU-trace driven mode executing traces of representative sections of our workloads collected with a custom Pintool [209]. We warm-up the caches by fast-forwarding 100 million (M) instructions. We simulate each representative trace for 500M instructions (for multi-core simulations, until each core executes at least 500M instructions).
We use the system configuration provided in Table 1 in our evaluations. Although our evaluation is based on DDR4 DRAM, the modifications required to enable SMD can be adopted in other DRAM standards, as we explain in section 6.
@@
| Memory Controller | 64-entry read/write request queue, FR-FCFS-Cap [160] with Cap=7 |
@@
DDR4 Baseline. The baseline system 1) uses per-rank refresh181818The DDR4 standard does not support per-bank refresh. and 2) does not perform ECC scrubbing.
@@
Comparison Points. We compare SMD to memory controller/DRAM co-design techniques. First, DARP [19] intelligently schedules per-bank refresh commands to idle banks while the memory controller is in write mode. Doing so reduces delays imposed by refresh operations on memory demand requests (e.g., load instructions executed by the processor). Second, DSARP [19] implements DARP and also modifies DRAM chip design and the DRAM interface to concurrently perform a refresh operation in one subarray and a memory access in another subarray, thereby hiding the latency of refresh operations and reducing delays they impose on memory requests. We modify DARP and DSARP such that they also reduce delays imposed on memory requests by Deterministic RowHammer Protection’s (section 5.2) victim row refresh operations.
@@
Parameter Values of Maintenance Mechanisms and SMD. SMD-FR (section 5.1) refreshes a DRAM row every . SMD-FR refreshes 8 DRAM rows when it locks a DRAM region (RG = 8) with each refresh operation and postpones up to 8 such refresh operations (N = 8). Based on [20], for SMD-VR, we conservatively assume 0.1% of rows in each bank need to be refreshed every while the rest retain their data correctly for and more. SMD-VR uses a 8K-bit Bloom Filter with 6 hash functions. SMD-PRP refreshes the victims of an activated row with a high probability, i.e., = 1%. SMD-DRP (section 5.2) refreshes the victims before the aggressor is activated (512) times during a tREFW (). SMD-MS (section 5.3) operates with an aggressive 5-minute scrubbing period.191919We analyze SMD-DRP, and SMD-MS at various configurations. Their performance overheads are relatively low even at more aggressive settings. We configure SMD with 16 lock regions (section 4.2) in a DRAM bank unless stated otherwise. SMD uses an ACT Retry Interval () of (section 4.3).
@@
Evaluated System Configurations. We evaluate 10 different systems: 1) the DDR4 baseline system as described in Table 1, 2) SMD-FR that can concurrently perform a refresh in a lock region and a memory access in another, 3) SMD-VR inherits SMD-FR’s abilities and refreshes different rows at different refresh rates, 4) SMD-FR + SMD-PRP inherits SMD-FR’s abilities and can concurrently perform a probabilistic victim row refresh operation in a lock region and a memory access in another, 5) SMD-FR + SMD-DRP inherits SMD-FR’s abilities and can concurrently perform a deterministic (based on Graphene’s [61] tracking algorithm) victim row refresh operation in a lock region and a memory access in another, 6) SMD-FR + SMD-MS202020We evaluate SMD-DRP and SMD-MS with SMD-FR as the refresh mechanism since SMD-FR is very simple and easy to implement. inherits SMD-FR’s abilities and performs in-DRAM memory scrubbing, 7) SMD-Combined implements all three SMD-based maintenance mechanisms (SMD-FR + SMD-DRP + SMD-MS), 8) DARP-Combined inherits the abilities of DARP [19] and implements the memory-controller-based versions of fixed rate refresh (FR), deterministic RowHammer protection (DRP), and memory scrubbing (MS), 9) DSARP-Combined inherits the abilities of DSARP [19] and implements the memory-controller-based (MC-based) versions of all three maintenance mechanisms, and 10) No-Refresh is a hypothetical system configuration that does not perform any maintenance operations. MC-based implementations of SMD maintenance mechanisms are functionally equivalent to SMD-based ones (e.g., SMD-Combined and MC-Combined are configured to provide the same RowHammer prevention guarantees).
8 Performance and Energy Results
We evaluate the performance and energy efficiency of SMD-based maintenance mechanisms. Our results show that 1) substantial system performance and DRAM energy benefits accompany the system robustness improvements provided by SMD that performs periodic refresh, RowHammer protection, and memory scrubbing in DRAM, 2) performance and energy of SMD incorporating these techniques is very similar to those of a hypothetical system that does not perform any DRAM maintenance operations, and 3) SMD outperforms a state-of-the-art work (DSARP [19]) that can concurrently perform a maintenance operation in one DRAM subarray and a memory access in another.
@@ 8.1 Single-core Performance
Fig. 8 shows the speedup of the 22 single-core workloads that have at least 10 Last-Level Cache Misses per Kilo Instruction (LLC MPKI) with SMD-FR, SMD-VR, SMD-FR + SMD-PRP, SMD-FR + SMD-DRP, SMD-FR + SMD-MS, SMD-Combined, and No-Refresh, over the baseline system that only performs memory-controller-based periodic refresh operations. We make five major observations.
First, SMD-FR provides 4.8% average speedup across all workloads. Although SMD-FR and DDR4 have similar average latency to refresh a single DRAM row, SMD-FR outperforms baseline DDR4 refresh due to two main reasons: SMD-FR 1) enables concurrent access and refresh of two different lock regions within a DRAM bank and 2) completely eliminates the REF commands on the DRAM command bus, thereby reducing the command bus utilization. Second, SMD-VR provides a higher 5.5% average speedup across all workloads as it mitigates the overhead of periodic refresh operations by refreshing different rows at different refresh rates. Third, SMD-FR + SMD-DRP (SMD-FR + SMD-PRP) provides 4.8% (4.2%) average speedup across all workloads. We observe that for a RowHammer threshold of 512 row activations (section 5.2), SMD-DRP and SMD-PRP can almost completely overlap the latency of victim row refresh operations in one lock region with the latency of a memory access to another lock region. Fourth, SMD-FR + SMD-MS provides 5.0% average speedup across all workloads. SMD enables low-overhead in-DRAM memory scrubbing for a relatively high memory scrubbing rate (each DRAM cell gets scrubbed every 5 minutes). The scrubbing operations induce negligible performance overhead: The average speedup provided by SMD-FR + SMD-MS is within <0.1% of SMD-FR if we exclude cond-mat-2003.212121SMD-FR + SMD-MS provides a relatively high 23.9% speedup for cond-mat-2003. SMD-FR provides a smaller speedup than SMD-Combined only for cond-mat-2003. For the same workload, SMD-FR has 5.1% higher average memory access latency than SMD-Combined. Even though SMD-Combined rejects (ACT_NACKs) more activate commands (15622) than SMD-FR (14286) on average across all four DRAM channels, in SMD-Combined, load instructions at the head of the processor core’s reorder buffer commit earlier on average than in SMD-FR. Consequently, SMD-Combined is less frequently bottlenecked by reorder buffer stalls (15.5% of execution time) than SMD-FR (40.2% of execution time). We attribute SMD-FR’s relatively poor performance for cond-mat-2003 to reorder buffer stalls. We hypothesize that the duration and the frequency of ACT_NACKs in SMD-FR affect main memory request scheduling in a way that older requests in the reorder buffer experience higher memory latency. Fifth, SMD-Combined provides 5.0% average speedup across all workloads. SMD-Combined provides 84.7% of the speedup provided by No-Refresh on average across all evaluated single-core workloads, while also improving system robustness with RowHammer protection and memory scrubbing (which No-Refresh does not do).
@@ 8.2 Multi-core Performance
Fig. 9 shows weighted speedup (normalized to the weighted speedup of the DDR4 baseline) of 60 four-core workloads (20 per memory intensity level) with SMD-FR, SMD-VR, SMD-FR + SMD-PRP, SMD-FR + SMD-DRP, SMD-FR + SMD-MS, SMD-Combined, and No-Refresh. The white circles inside each box (error lines) represent the average (minimum and maximum) normalized weighted speedup across the 20 workloads in the corresponding group. We make three major observations.
First, SMD provides substantial performance improvements for memory intensive workloads with relatively high LLC MPKI (i.e., 4c-medium and 4c-high workloads). For example, SMD-Combined provides 5.1% (8.2%) and 8.9% (10.5%) average (maximum) speedup across all evaluated 4c-medium and 4c-high workloads, respectively. SMD does so by overlapping the latency of a majority of the maintenance operations with the latency of memory demand requests (e.g., load instructions). For 4c-high workloads, SMD-Combined provides 88.3% of the average speedup provided by No-Refresh even though SMD-Combined performs periodic refresh (section 5.1), RowHammer protection (section 5.2), and frequent memory scrubbing (section 5.3) that No-Refresh does not perform. Second, all evaluated SMD configurations provide similar speedups. For example, SMD-FR, SMD-FR + SMD-DRP, SMD-FR + SMD-MS, and SMD-Combined all provide approximately 8.9%; SMD-VR provides 9.2% and SMD-FR + SMD-PRP provides 7.4% average speedup across the evaluated 4c-high workloads. This indicates that SMD is able to effectively hide the latencies of RowHammer protection and memory scrubbing operations. Third, SMD provides small performance improvements for non-memory-intensive workloads with relatively low LLC MPKI (i.e., 4c-low workloads). For such workloads, completely eliminating maintenance operations (i.e., No-Refresh) also provides small performance improvements. SMD-Combined provides 88.8% of the speedup provided by No-Refresh for 4c-low workloads.
@@
We conclude that SMD provides substantial system performance benefits by concurrently performing a maintenance operation and one or more memory access in different lock regions in a bank.
8.3 Energy Consumption